OpenStudy (anonymous):

What is the solution of: Y=2x+7 X=-3

OpenStudy (marigirl):

what we can do to find the point where both graphs cross over is to substitute x=-3 into the equation y=2x+7 y=2(-3)+7 y=-6+7 y=1 so the point of intersection (or the point where the graphs cross over) is at (-3,1)